A friction trolley is pushed from rest with a force of 40N for 15s. It strikes a wall and comes to rest in 0.2s. Compute the average force that the wall exerted on the trolley.


1
Expert's answer
2021-02-03T16:12:48-0500

The momentum initially given to the trolley is:


"p = F_1t_1"

where "F_1 = 40N", and "t = 15s". The wall took all this momentum in "t_2 = 0.2s". Thus, the average force exerted on the trolley is:

"F_2 = \dfrac{p}{t_2} = \dfrac{F_1t_1}{t_2}\\
F_2 = \dfrac{40\cdot 15}{0.2} = 3000N"

Answer. 3000 N.